# GraphToxin: Reconstructing Full Unlearned Graphs from Graph Unlearning

### v1

@@ Abstract
Graph unlearning has emerged as a promising solution for complying with “the right to be forgotten” regulations by enabling the removal of sensitive information upon request. However, this solution is not foolproof. The involvement of multiple parties creates new attack surfaces, and residual traces of deleted data can still remain in the unlearned graph neural networks (GNNs). These vulnerabilities can be exploited by attackers to recover the supposedly erased samples, thereby undermining the inherent functionality of graph unlearning.
In this work, we propose GraphToxin, the first graph reconstruction attack against graph unlearning. Specifically, we introduce a novel curvature matching module to provide a fine-grained guidance for full unlearned graph recovery. We demonstrate that GraphToxin can successfully subvert the regulatory guarantees expected from graph unlearning—it can recover not only a deleted individual’s information and personal links but also sensitive content from their connections, thereby posing substantially more detrimental threats. Furthermore, we extend GraphToxin to multiple node removal under both white-box and black-box setting, showcasing its practical feasibility and its potential to cause considerable harm.
We highlight the necessity of a worst-case analysis and propose a comprehensive evaluation framework to systematically assess the attack performance under both random and worst-case node removal scenarios. This provides a more robust and realistic measure of the vulnerability of graph unlearning methods to graph reconstruction attacks. Our extensive experiments demonstrate the effectiveness and flexibility of GraphToxin. Notably, we show that existing defense mechanisms are largely ineffective against this attack and, in some cases, can even amplify its performance. Given the severe privacy risks posed by GraphToxin, our work underscores the urgent need for the development of more effective and robust defense strategies against this type of attacks.
@@ 1 Introduction
Graph neural networks (GNNs) have demonstrated remarkable success across diverse domains, including social network analysis, drug discovery, and financial risk management. Their strength lies in capturing complex structural dependencies through iterative message passing. However, this mechanism inherently amplifies the risk of sensitive information leakage, raising significant user privacy concerns. Concurrently, data protection regulations such as the General Data Protection Regulation (GDPR) [gdpr] enforce the “right to be forgotten”, requiring GNNs’ capacity to remove personal data upon request. These privacy and legal imperatives have spurred the rapid emergence of graph unlearning (GU) [graph_unlearn].
@@
Graph unlearning aims to scrub the influence of specific nodes, edges, or subgraphs from a trained GNN while preserving model performance on the remaining graph. Specifically, upon an unlearning request, the model owner typically employs parameter adjustments or gradient-based optimization to mitigate these effects instead of full retraining—which is prohibitively expensive and time-consuming—and subsequently releases the unlearned version. Despite its success, GU inevitably leaves residual traces of the “forgotten” targets. Although such remnants are often small enough to avoid noticeable degradation in overall performance, they can still be exploited by adversaries through membership inference [unlearning_inversion_attacks_graph] or graph reconstruction attacks, thereby undermining privacy guarantees. Moreover, the involvement of multiple parties further expands the attack surface. For instance, the service providers or platforms can retain copies of the previous models prior to unlearning while users can download such models locally on their devices. These decentralized behaviors fundamentally undermine the assurance of GU. In this work, we focus on the most challenging scenario—node unlearning [surveygraphunlearning] for node-level tasks—and present the first full graph reconstruction attack against graph unlearning (GU-FGRA).
@@
Research Gap—Graph Unlearning To the best of our knowledge, only one prior work [unlearning_inversion_attacks_graph] has investigated the privacy risks of GU. However, their study is limited to a single edge removal and only aims to infer the link membership of a given node pair, rather than reconstructing the full deleted graph topology. Consequently, their approach solely reveals minimal information and fails to recover confidential or sensitive content encoded in node features. Additionally, their attack assumptions—access to partial unlearned/remaining graphs and a shadow graph with a distribution similar to that of the original graph—are unrealistic in practice due to strict security and privacy policies. Even if partial data were leaked, finding a faithful replica of the original dataset is infeasible and using publicly available data from a different distribution severely degrades the quality of reconstructed graphs due to distributional shifts.
@@
Graph Reconstruction Attack (GRA) Regarding GRA, most existing studies primarily focus on inferring link memberships [link_steal_gnn, link_teller]. Only a few attempt to recover sensitive node features [mia_gra_gnn, aia_gnn] or reconstruct the entire graph topology [inference_gnn, explain_gra]. This inconsistent terminology can be misleading, as a true full GRA should reveal all graph components. Moreover, current research typically rely on access to node embeddings or shadow models, and often assume that sensitive node features or portions of the topology are available for recovery. They also lack comprehensive evaluation, hindering their practical applicability in scenarios such as red-teaming and model auditing.
@@
Machine Unlearning Existing inversion attacks against MU cannot be trivially adapted to GU-FGRA, as the complex topological structures and rich node features in graphs produce entangled patterns during GNN message passing. To date, only one work [first_unlearn_attack] has explored single sample reconstruction in MU. However, it is limited to partial recovery of deleted samples under a white-box setting and performs poorly when targeting multiple deletions. Under the i.i.d. assumption, its multiple removals can be treated as a series of independent single removals. In contrast, graph unlearning is fundamentally different: due to the “ripple effects” propagated through GNNs, even a single node removal can affect its neighbors and higher-order structures, preventing a straightforward extension to GU.
@@
    Is it possible to recover full unlearned graphs under realistic white-box settings and how this attack can be extended to multiple node removals?
@@
    Can this attack be generalized to the black-box setting and reconstruct full unlearned graphs using only minimal prior knowledge for multiple node removals?
@@
    How can we rigorously and comprehensively evaluate the performance of this full reconstruction attack?
@@
Unique Challenges and Strategies Beyond the severe challenges previously discussed, the feasibility of gradient-based GU-FGRA under realistic attack settings remains underexplored. Since no shadow graphs or models are available and only minimal prior knowledge is mastered, even a simple unlearning attack targeting a single node is an intricate task. Notably, our goal is to reveal full unlearned graphs rather than a simple combination of multiple attribute inference attacks of the deleted node and multiple link membership inference attacks of its $1$-hop neighbors. This challenge is further exacerbated for multiple node removals due to much worse structural and feature entanglements and ultimately more obscure gradient proxy for full graph recovery.
@@
To answer the key research questions and to address their unique challenges under a realistic white-box setting, we present GraphToxin, the first full graph reconstruction attack against graph unlearning. Since the difference of publicly leaked gradients from the original GNN and its unlearned GNN implicitly encodes the information of the deleted graph, GraphToxin exploits this gradient difference for full graph recovery through three core modules. First, the gradient matching module acts as a holistic roadmap for full unlearned graph recovery by bridging the gap between the ground-truth gradient difference and gradients from the generated unlearned graph, however, it only yields a coarse proxy. Second, a novel curvature matching module provides fine-grained signals by penalizing the directions of high curvature to ultimately produce a more plausible recovery solution. Finally, a feature smoothness module is utilized to capture and enforce similarity among connected nodes.
@@
GraphToxin also tackles the intractable challenge of the data-free black-box setting. This scenario presents a major hurdle since attackers are restricted to using the model’s posterior probabilities only instead of its gradients. This limitation inevitably introduces approximation errors, as black-box GU-FGRA must attempt to infer fine-grained graph details from coarse posterior outputs, which can significantly degrade the reconstruction quality. To mitigate this issue, GraphToxin introduces a semantic calibration module. The goal of this module is to minimize the discrepancy between the ground-truth labels and the predictions for the recovered nodes, thereby improving the attack’s fidelity.
@@
The efficacy of GraphToxin is validated through a rigorous and comprehensive evaluation framework. We contend that the conventional setting of our attack target is unconvincing as it focuses solely on small-scale random removals of the training graph [graph_unlearn, GIF, MEGU, ETR]. Given the inherent sparsity of real-world graphs, this operation often selects nodes that are less influential for model predictions. These “unimportant” nodes typically possess low degrees and cause negligible influences after removal. However, in reality, attackers would prioritize high-value targets. To better mimic actual attacks, we propose a worst-case unlearning analysis. In this research, we define this worst-case scenario as the removal of the top $10\%$ of nodes with the highest degrees. Furthermore, we argue that existing evaluation metrics are often simplistic and unreliable [gradient_attack_gnn, graph_mia, grain]. Accordingly, we develop a systematic evaluation framework that incorporates reconstruction-related metrics to jointly measure both recovered features and topology from a global view, and performance-level metrics to assess the functional utility of the full recovered graph.
@@
    Pioneering Attack: We design GraphToxin, the first full graph reconstruction attack against graph unlearning. It not only recovers the deleted nodes and their graph topology, but also steals the sensitive information of their neighbors, posing more detrimental privacy risks.
@@
    Scalability and Generalization: We successfully extend GraphToxin to multiple node removals with comparable attack performance. We also generalize GraphToxin to the black-box setting with only minimal prior knowledge.
@@
    Rigorous Evaluation Framework Design: We propose a rigorous evaluation framework grounded in feature-level, global-level, and performance-level perspectives. We also investigate GraphToxin after removing the most influential nodes to assess its attack potency in reality.
@@
    Empirical Validation: We conduct extensive experiments across diverse graph datasets, GNN backbones, and both exact and approximate graph unlearning methods. The results demonstrate the effectiveness and flexibility of GraphToxin in both white-box and black-box settings, for a single or multiple node removals.
@@
    Highlighting Urgency for Robust Defenses: We empirically show that current defense methods fail to sufficiently mitigate or even enhance the attack performance of GraphToxin. This finding highlights the urgent need for developing more effective defense mechanisms against this potent attack.
@@ 2.1 Notations
Given an undirected attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$, $\mathcal{V}$ denotes the node set with cardinality $|\mathcal{V}|$ and each node $v$ is associated with a feature vector $X_{v}\in\mathbb{R}^{1\times\mathcal{D}}$ and a label vector $Y_{v}\in\mathbb{R}^{1\times|Y|}$, where $\mathcal{D}$ is the feature dimension and $|Y|$ denotes the number of labels. $\mathcal{E}$ denotes the set of edges with $|\mathcal{E}|$ elements, which can be represented by an adjacency matrix $A\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $A_{u,v}=1$ iff $(u,v)\in\mathcal{E}$, and $A_{u,v}=0$ otherwise. For simplicity, we denote the graph as $\mathcal{G}=(A,X)$. $\mathcal{F}_{ori}$, $\mathcal{F}_{un}$ and $\mathcal{F}_{att}$ indicate the original GNN, unlearned GNN, and GU-FGRA, respectively.
@@ 2.3 Graph Unlearning
We only consider the most challenging GU—node unlearning throughout our paper, which is formally defined as:
@@
Given the original training graph $\mathcal{G}=(A,X)$, the original trained GNN $\mathcal{F}_{ori}$ and a specific unlearning request $\Delta\mathcal{G}_{v}=(v,\mathcal{E}_{\mathcal{N}^{1}_{v}},X_{v})$, node unlearning aims to unlearn a single node $v$, its feature $X_{v}$ and its links to 1-hop neighbors $\mathcal{N}^{1}_{v}$ while minimizing the discrepancy between the unlearned model $\mathcal{F}_{{un}_{\Delta\mathcal{G}_{v}}}$ and the fully retrained model $\mathcal{F}_{\mathcal{G}\backslash\Delta\mathcal{G}_{v}}$.
@@
Upon multiple unlearning requests $\Delta\mathcal{G}=(\mathcal{V}_{d},\mathcal{E}_{d},X_{d})$, node unlearning aims to unlearn multiple nodes $\mathcal{V}_{d}$, their node features $X_{d}$ and links to their 1-hop neighbors $\mathcal{E}_{d}$, while minimizing the discrepancy between the unlearned model $\mathcal{F}_{{un}_{\Delta\mathcal{G}}}$ and the fully retrained model $\mathcal{F}_{\mathcal{G}\backslash\Delta\mathcal{G}}$. Different from single node unlearning, $\mathcal{E}_{d}$ is much more complicated rather than a simple combination of multiple $\mathcal{N}^{1}_{v},v\in\mathcal{V}_{d}$.
@@ 2.4 Graph Reconstruction Attack (GRA)
Given a trained GNN $\mathcal{F}$ and attack knowledge $\mathcal{K}$, GRA $\mathcal{F}^{-1}$ endeavors to recover sensitive information $\mathcal{G}_{par}$ of its training graph $\mathcal{G}$ by minimizing the distance between the reconstructed component $\hat{\mathcal{G}}_{par}$ and its ground-truth $\mathcal{G}_{par}$, where $\hat{\mathcal{G}}_{par}=\mathcal{F}^{-1}(\mathcal{K})$ and $\mathcal{G}_{par}\in\mathcal{G}$.
Based on the type of $\mathcal{K}$, general GRA can be specialized to white-box GRA when $\mathcal{K}$ contains model gradients $\nabla\mathcal{L}(\theta)$ (where $\theta$ represents model parameters); it transforms into gray-box GRA when $\mathcal{K}$ covers node embeddings but no white-box knowledge; and it presents black-box GRA when none of the above information is available.
Based on the recovered information $\hat{\mathcal{G}}_{par}$, general GRA can target specific components: recovering partial or full node features through attribute inference attacks, inferring a specific link membership via link stealing attacks, or reconstructing topological structures through graph topology reconstruction attacks. Our research tackles the most challenging objective—fully revealing node features and graph topology of unlearned graphs through full GRA. We provide a clear comparison of these attacks in Table I.
TABLE I: Attack Comparison. $\surd$ represents “access” or “applicable”, $\times$ represents “no access”, $\star$ denotes “shadow graph”, and $\varnothing$ denotes “not applicable” or “unknown”. N/A means unexplored in the field.
Attack Type Attack Knowledge $\mathcal{K}$ Recovery Content $\mathcal{G}_{par}$ Attack Unlearning Case
@@
Attribute Inference Attack (N/A) $\surd\star$ Grey/Black Box $\varnothing$ Single/Partial $\varnothing$ $\varnothing$ $\varnothing$
Link Stealing Attack [unlearning_inversion_attacks_graph] $\surd\star$ Grey/Black Box Partial Unlearned Graph $\varnothing$ Edge $\surd$ $\varnothing$
Graph Topology Reconstruction Attack (N/A) $\surd\star$ Grey/Black Box $\varnothing$ $\varnothing$ Partial/Full $\varnothing$ $\varnothing$
Full Graph Reconstruction Attack (Ours) $\times$ White/Grey/Black Box Simple Unlearning Statistics Full Full $\surd$ $\surd$

### v2

@@ Abstract
Graph unlearning has emerged as a promising solution to comply with “the right to be forgotten” regulations by enabling the removal of sensitive information upon request. However, this solution is not foolproof. The involvement of multiple parties creates new attack surfaces, and residual traces of deleted data can still remain in the unlearned graph neural networks (GNNs). These vulnerabilities can be exploited by attackers to recover the supposedly erased samples, thereby undermining the intended functionality of graph unlearning.
In this work, we propose GraphToxin, the first full graph reconstruction attack against graph unlearning. Specifically, we introduce a novel curvature matching module to provide fine-grained guidance for unlearned graph recovery. We demonstrate that GraphToxin can successfully subvert the regulatory guarantees expected from graph unlearning—it can recover not only a deleted individual’s information and personal links but also sensitive content from their connections, thereby posing substantially more detrimental threats. Furthermore, we extend GraphToxin to multiple-node removal under both white-box and black-box settings, showcasing its practical feasibility and potential to cause considerable harm.
We highlight the necessity of worst-case analysis and propose a systematic evaluation framework to assess attack performance under both random and worst-case node removal scenarios. Our extensive experiments demonstrate the effectiveness and flexibility of GraphToxin. Notably, existing defense mechanisms are largely ineffective against this attack or even amplify its performance in some cases. Given the severe privacy risks posed by GraphToxin, our work underscores the urgent need for more effective and robust defenses.
@@ 1 Introduction
Graph neural networks (GNNs) have demonstrated remarkable success across diverse domains, including social network analysis, drug discovery, and financial risk management [gnn_app]. Their strength lies in capturing complex structural dependencies through iterative message passing. However, this mechanism inherently amplifies the risk of sensitive information leakage, raising significant user privacy concerns. Concurrently, data protection regulations such as the General Data Protection Regulation (GDPR) [gdpr] enforce the “right to be forgotten”, requiring GNNs’ to remove personal data upon request. These privacy and legal imperatives have spurred the rapid emergence of graph unlearning (GU) [graph_unlearn].
@@
Graph unlearning aims to remove the influence of specific nodes, edges, or subgraphs from a trained GNN while preserving model performance on the remaining graph. Specifically, upon receiving an unlearning request, the model owner typically employs parameter adjustments or gradient-based optimization to mitigate these effects instead of full retraining, which is prohibitively expensive and time-consuming, and subsequently releases the unlearned version. Despite its success, GU inevitably leaves residual traces of the “forgotten” targets. Although such remnants are often small enough to avoid noticeable performance degradation, they can still be exploited by adversaries through membership inference [unlearning_inversion_attacks_graph] or graph reconstruction attacks, thereby undermining privacy guarantees. Moreover, the involvement of multiple parties further expands the attack surface. For instance, the service providers or platforms can retain copies of the previous models prior to unlearning for legitimate reasons such as backup and debugging, while users can download such models locally on their devices. These decentralized behaviors fundamentally undermine the assurance of GU. In this work, we focus on the most challenging scenario—node unlearning [surveygraphunlearning] for node-level tasks—and present the first full graph reconstruction attack against graph unlearning (GU-FGRA).
@@
Research Gap—Graph Unlearning To the best of our knowledge, only one prior work [unlearning_inversion_attacks_graph] has investigated the privacy risks of GU. However, their study is limited to single-edge removal and aims to infer the link membership of a given node pair, rather than reconstructing the full unlearned graph topology. Consequently, their approach solely reveals minimal information and fails to recover confidential or sensitive content encoded in node features. Additionally, they assume that the adversary has access to partial unlearned/remaining graphs and a shadow graph with a distribution similar to that of the original graph, which are unrealistic in practice due to strict security and privacy policies. Even if partial data were leaked, finding a faithful replica of the original dataset is infeasible and using publicly available data from a different distribution severely degrades the quality of reconstructed graphs due to distributional shifts.
@@
Graph Reconstruction Attack (GRA) With respect to GRA, existing studies primarily focus on inferring link memberships [link_steal_gnn, link_teller]. Only a few attempt to recover sensitive attributes [mia_gra_gnn, aia_gnn] or the full graph topology [inference_gnn, explain_gra]. This inconsistent terminology can be misleading, as a true full GRA should reveal all graph components. Moreover, current research typically rely on access to node embeddings or shadow models, and often assume that sensitive node features or portions of the topology are available for recovery. They also lack comprehensive evaluation, hindering their practical applicability in scenarios such as red-teaming and model auditing.
@@
Machine Unlearning (MU) Existing reconstruction attacks against MU cannot be trivially adapted to GU-FGRA, as the complex topological structures and rich node features in graphs produce entangled patterns during GNN message passing. To date, a few works have explored reconstruction in MU. These efforts are either limited to linear models for exact unlearning [bertran2024reconstruction] or recover only partial removed samples under white-box settings, with poor scalability to multiple-node removal [first_unlearn_attack]. Under the i.i.d. assumption commonly adopted for tabular data, multiple sample removals can be treated as a sequence of independent single removals. In contrast, graph unlearning is fundamentally different: due to the “ripple effects” propagated through GNNs, even a single node removal can affect its neighbors and higher-order structures, preventing a straightforward extension to GU.
@@
    Is it possible to recover full unlearned graphs under realistic white-box settings and how this attack can be extended to multiple-node removal?
@@
    Can this attack be generalized to black-box settings and reconstruct full unlearned graphs using only minimal prior knowledge for multiple-node removal?
@@
    How can we comprehensively evaluate the performance of this full reconstruction attack?
@@
Unique Challenges and Strategies Beyond the challenges previously discussed, the feasibility of gradient-based GU-FGRA under realistic attack settings remains underexplored. Since no shadow graphs or models are available and only minimal prior knowledge is mastered, even a simple unlearning attack targeting a single node is an intricate task. Notably, our goal is to reveal full unlearned graphs rather than a simple combination of multiple attribute inference attacks of the removed node and multiple link membership inference attacks of its $1$-hop neighbors. This challenge is further exacerbated under multiple-node removal, where more complex structural and feature entanglements yield increasingly ambiguous gradient proxies for full graph recovery.
@@
To answer the key research questions and to address their unique challenges under realistic white-box settings, we present GraphToxin, the first full graph reconstruction attack against graph unlearning. Since the difference between publicly accessible gradients of the original and unlearned GNNs implicitly captures the information associated with the unlearned graph, GraphToxin exploits this difference for full graph recovery through three core modules. First, a gradient matching module acts as a holistic roadmap for full unlearned graph recovery by bridging the gap between the ground-truth gradient difference and the gradients generated from the recovered unlearned graph, however, it only yields a coarse proxy. Second, a novel curvature matching module provides fine-grained signals by downweighting the directions of high curvature to ultimately produce a more plausible recovery solution. Finally, a feature smoothness module is utilized to enforce similarity among connected nodes.
@@
GraphToxin also tackles the intractable challenge of the data-free black-box setting. This scenario presents a major hurdle since the adversary is restricted to using only the model’s posterior probabilities instead of its gradients. This limitation inevitably introduces approximation errors, as black-box GU-FGRA must attempt to infer fine-grained graph details from coarse posterior outputs, which can degrade the reconstruction quality. To mitigate this issue, GraphToxin introduces a semantic calibration module. The goal of this module is to minimize the discrepancy between the ground-truth labels and the predictions for the recovered nodes, thereby encouraging semantically consistent reconstructions.
@@
The effectiveness of GraphToxin is evaluated through a comprehensive evaluation framework. Existing GU methods primarily focus on small-scale random removals, which provides an incomplete assessment of reconstruction risks. [graph_unlearn, GIF, MEGU, ETR]. Given the inherent sparsity of real-world graphs, this operation often selects nodes that are less influential for model predictions. These “unimportant” nodes typically possess low degrees and cause negligible influences after removal. However, in reality, attackers would prioritize high-value targets. To better mimic actual attacks, we propose a worst-case evaluation. In our work, we define this worst-case scenario as the removal of the top $10\%$ of nodes with the highest degrees. Furthermore, traditional evaluation metrics are often insufficient to characterize reconstruction quality [gradient_attack_gnn, graph_mia, grain]. Motivated by this limitation, we develop a systematic evaluation framework that incorporates reconstruction-related metrics to capture the global consistency of recovered features and topology, and performance-level metrics to assess functional utility of the full recovered graph.
@@
    Pioneering Attack: We design GraphToxin, the first full graph reconstruction attack against graph unlearning. It not only recovers the removed nodes and their graph topology, but also steals the sensitive information of their neighbors, posing more detrimental privacy risks.
@@
    Scalability and Generalization: We successfully extend GraphToxin to multiple-node removal with comparable attack performance, and generalize GraphToxin to the black-box setting with only minimal prior knowledge.
@@
    Rigorous Evaluation Framework Design: We propose a comprehensive evaluation framework grounded in feature-level, global-level, and performance-level perspectives. Additionally, we assess GraphToxin after removing the most influential nodes to assess its attack potency in reality.
@@
    Empirical Validation: We conduct extensive experiments across diverse graph datasets, GNN backbones, and both exact and approximate graph unlearning methods. The results demonstrate the effectiveness and flexibility of GraphToxin in both white-box and black-box settings, for single or multiple-node removal.
@@
    Highlighting Urgency for Robust Defenses: We empirically show that existing defense methods are ineffective against GraphToxin and in some cases may even amplify its attack performance. This finding highlights the urgent need for developing more effective defense mechanisms against this potent attack.
@@ 2.1 Notations
Given an undirected attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$, $\mathcal{V}$ denotes the set of nodes with $|\mathcal{V}|$ elements and each node $v$ is associated with a feature vector $X_{v}\in\mathbb{R}^{1\times d}$ and a label vector $Y_{v}\in\mathbb{R}^{1\times|Y|}$, where $d$ is the feature dimensionality and $|Y|$ denotes the number of labels. $\mathcal{E}$ denotes the set of edges with $|\mathcal{E}|$ elements, which can be represented by an adjacency matrix $A\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $A_{w,v}=1$ iff $(w,v)\in\mathcal{E}$, and $A_{w,v}=0$ otherwise. For simplicity, we denote the graph as $\mathcal{G}=(A,X)$. $\mathcal{F}_{o}$, $\mathcal{F}_{u}$ and $\mathcal{F}_{a}$ indicate the original GNN, unlearned GNN, and attack model, respectively.
@@ 2.3 Graph Unlearning
We consider the most challenging GU—node unlearning throughout our paper, which is formally defined as:
@@
Given the original training graph $\mathcal{G}=(A,X)$, the original trained GNN $\mathcal{F}_{o}$ and a single unlearning request $\Delta\mathcal{G}_{v}=(v,\mathcal{E}_{\mathcal{N}^{1}_{v}},X_{v})$, node unlearning aims to unlearn a single node $v$, its feature $X_{v}$ and its links to 1-hop neighbors $\mathcal{N}^{1}_{v}$ by training an unlearned model $\mathcal{F}_{u}$ on the remaining graph $\mathcal{G}\backslash\Delta\mathcal{G}_{v}$ that approximates the fully retrained model $\mathcal{F}_{o}^{\prime}(\mathcal{G}\backslash\Delta\mathcal{G}_{v})$.
@@
Upon receiving multiple unlearning requests $\Delta\mathcal{G}=(\mathcal{V}_{\mathcal{M}},\mathcal{N}^{1}_{\mathcal{M}},X_{\mathcal{M}})$, node unlearning aims to unlearn multiple nodes $\mathcal{M}$, their node features $X_{\mathcal{M}}$ and links to their 1-hop neighbors $\mathcal{N}^{1}_{\mathcal{M}}$ by training unlearned model $\mathcal{F}_{u}$ on the remaining graph $\mathcal{G}\backslash\Delta\mathcal{G}$ that approximates the fully retrained model $\mathcal{F}_{o}^{\prime}(\mathcal{G}\backslash\Delta\mathcal{G})$.
@@ 2.4 Graph Reconstruction Attack (GRA)
Given access to a GNN $\mathcal{F}_{o}$ trained on graph $\mathcal{G}$, a GRA $\mathcal{F}_{a}$ endeavors to recover sensitive information $\hat{\mathcal{G}}_{par}$ that is indistinguishable from its ground-truth information $\mathcal{G}_{par}\in\mathcal{G}$.
Depending on the adversary’s interaction with $\mathcal{F}_{o}$, a general GRA reduces to a white-box setting when the adversary can access model gradients $\nabla\mathcal{L}(\theta)$ (hereinafter abbreviated as $\nabla\mathcal{L}$), where $\theta$ denotes model parameters. Under the gray-box setting, GRA allows the adversary to access intermediate model outputs such as node embeddings but no gradient-level information, whereas the black-box setting restricts the adversary to only query the victim model, without access to internal representations or gradients.
Based on the recovered information $\hat{\mathcal{G}}_{par}$, a general GRA can be instantiated to target different graph components. Specifically, it can recover partial or complete node features via attribute inference attacks, infer the existence of specific edges through link stealing attacks, or reconstruct structural information via graph topology reconstruction attacks. In this work, we focus on the most challenging setting, namely full GRA, which aims to jointly recover both full node features and graph topology of the target graph. A systematic comparison of these attack paradigms is provided in Table 1.
Table 1: Attack Comparison. $\surd$ represents “access" or “applicable", and $\varnothing$ denotes “not applicable" or “unknown". N/A means currently unexplored in the field.
Attack Type Attacker’s Knowledge and Capabilities Recovery Content Attack Unlearning Case
@@
Attribute Inference Attack (N/A) Partial/Shadow Graph Grey/Black Box $\varnothing$ Single/Partial $\varnothing$ $\varnothing$ $\varnothing$
Link Stealing Attack [unlearning_inversion_attacks_graph] Partial/Shadow Graph Grey/Black Box Partial Unlearned Graph $\varnothing$ Single Edge $\surd$ $\varnothing$
Graph Topology Reconstruction Attack (N/A) Partial/Shadow Graph Grey/Black Box $\varnothing$ $\varnothing$ Partial/Full $\varnothing$ $\varnothing$
Full Graph Reconstruction Attack (Ours) Data-free White/Black Box #of Node/Edge/Label Changes Full Full $\surd$ $\surd$